OpenStudy (anonymous):

how do i solve (2^x-3)^2+2=9

OpenStudy (anonymous):

idk tell me the answer

OpenStudy (anonymous):

1. subtract 2 on both sides 2. then take square root on both sides. 3. add 3 to both sides. 4. take log on both sides

OpenStudy (anonymous):

log(3+/-sqrt(7),2)

OpenStudy (anonymous):

how do i take the square root on both sides. would that be? Square root of 2^x-3= square root of 7

OpenStudy (anonymous):

that is right.

OpenStudy (anonymous):

then i have square root of 2^x = square root of 7+3, if that is correct. how do i do the log

OpenStudy (anonymous):

no. you have \[2^x= \sqrt 7 +3\]

OpenStudy (anonymous):

log(2^x) = log(sqrt(7) + 3) x = log(sqrt(6) + 3)/log(2)

OpenStudy (anonymous):

Okay, i see. now how do i explain the log

OpenStudy (anonymous):

what do you mean? logging something is the opposite of exponentiating something

OpenStudy (anonymous):

Yes, i know that much. I'm sorry, let me rephrase that. What are the steps you took to get that answer for log. I can't just put an answer because that does me no good.

OpenStudy (anonymous):

so far i have log(2^x) =log (root(7)+3)

OpenStudy (anonymous):

log a^b = b log a so log 2^x = x log 2 = log(sqrt7 +3) x = log(sqrt7 +3)/log 2

OpenStudy (anonymous):

Honestly, I am lost after x log 2. I keep getting an end result of x=log2 (sq root(7) +3). Can you try and explain to me why I am wrong. Maybe that might help

OpenStudy (anonymous):

(2^x\[(2^x - 3 )^2 = 7 \rightarrow 2^{2x} - 3 \times 2^{x+1} + 9 = 7\] let \[2^x = y\] so \[y^2 - 3 \times 2 \times y + 2 = 0\] -> Solving.... y` = 3 + sqrt{7} `n y``= 3 - sqrt{7} back to 2^x = y .... \[x = \log_{2} {3\pm \sqrt{7}}\]